OpenStudy (anonymous):

use the distributive number to multiply. c (d + 11)

OpenStudy (anonymous):

*property NOT number

hero (hero):

General Distributive Rule: a(b + c) = ab + ac Applied Disbributive Rule c(d + 11) = cd + 11c

OpenStudy (anonymous):

You can think of it like this: c is multiplied by both d and 11, so c times d AND c times ll. c(d + 11) = cd + 11c

OpenStudy (anonymous):

thanks